OpenStudy (anonymous):

1) Draw a card 8 times from a shuffled deck, with replacement. What is the probability that at least 5 of the draws will be face cards? 2) Roll a six-sided die 8 times. What is the probability that 6 will appear at least 5 times? 3) The probability that a random newborn is male is roughly 0.51 . What is the probability that at least 2 of 4 newborns will be male?

OpenStudy (anonymous):

at least 5 times means 5, 6, 7, or 8 have to compute them all, then add

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

if you put \(x\) as the number of face cards drawn in 8 tries, do you know how to compute \[P(x=5)\]?

OpenStudy (anonymous):

note i need to some hint

OpenStudy (anonymous):

binomial

OpenStudy (anonymous):

\[P(x=k)=\binom{n}{k}p^k(1-p)^{n-k}\] is the formula

OpenStudy (anonymous):

k

OpenStudy (anonymous):

n= 8 and k= 5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

p=?

OpenStudy (anonymous):

you know how many face cards in a deck?

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

12 is right, three of each suit

OpenStudy (anonymous):

j ,k q

OpenStudy (anonymous):

right and how many cards in the deck?

OpenStudy (anonymous):

52

OpenStudy (anonymous):

so \(p=\)?

OpenStudy (anonymous):

p= 52?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

p is a probability has to be a number between zero and one there are 13 face cards and 52 cards total, so what is the probability you draw a face card?

OpenStudy (anonymous):

13 face card?

OpenStudy (anonymous):

draw the card 8 time

OpenStudy (anonymous):

i sense you are lost

OpenStudy (anonymous):

the p in the binomial distribution is the probability you get a success on any ONE trial

OpenStudy (anonymous):

k

OpenStudy (anonymous):

in this case p is the probability you draw a face card on one pick

OpenStudy (anonymous):

from those pick 5 of those will be face. that wat it mean ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

p is the probability you draw a face card on one pick

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i don't know another way to say it

OpenStudy (anonymous):

pick a card what is the probability it is a face card?

OpenStudy (anonymous):

1/13

OpenStudy (anonymous):

no there are 3 per suit, so \(\frac{3}{13}\)

OpenStudy (anonymous):

\[\left(\begin{matrix}8 \\ 5\end{matrix}\right)\frac{ 3 }{ 13 }^5(1-\frac{ 3 }{ 5 })^3\]

OpenStudy (anonymous):

3/13 mistype

OpenStudy (anonymous):

\[56 \frac{ 243 }{ 371293 }(\frac{ 1000 }{ 2197 })\]

OpenStudy (anonymous):

somehow i got 0.01668, I may be wrong